OpenStudy (anonymous):

c^8 d^-1 --------- c^-4 d^-8 how would I do this?

OpenStudy (astrophysics):

Use the following \[d^{-1} \implies \frac{ 1 }{ d },~~~c^{-4} = c^4\] Also when you're dividing exponents you subtract.

OpenStudy (astrophysics):

That should be \[\frac{ 1 }{ c^{-4} } = c^4\]

OpenStudy (astrophysics):

 or you can do it like this, lots of ways, with exponents :d